on average 3 traffic accidents per month occur at a certain intersection. what is the probability that in any given month at thisintersection exactly 5 accidents will occur?
1
Expert's answer
2020-12-10T10:52:35-0500

Here we have to make an assumption that number of accidents at a certain intersection follows Poisson distribution. Let "X" be the random

variable denoting the number of accidents in a month at a particular intersection:

"X\sim Po(\lambda)"


"P(X=x)=\dfrac{e^{-\lambda}(\lambda)^x}{x!}"

Given "\lambda=3"

The probability that in any given month at this intersection exactly 5 accidents will occur is


"P(X=5)=\dfrac{e^{-3}(3)^5}{5!}=\dfrac{81}{40}e^{-3}\approx0.100819"